OpenStudy (anonymous):

I need help to find the lcm of 12, 28, it is taking to long and i never get the right answer!

OpenStudy (anonymous):

Becca 12 written as prime numbers is 2 * 2 * 3 do you understand this?

OpenStudy (anonymous):

28 is 2 * 2 * 7

OpenStudy (anonymous):

Uhhh not really....

OpenStudy (anonymous):

How about this... 12 is 4 * 3 28 is 4 * 7

OpenStudy (anonymous):

The LCM of 12 and 28 is 84.

OpenStudy (anonymous):

Yea :V

OpenStudy (anonymous):

I get that blexing

OpenStudy (anonymous):

OK.. they both have a 4 so you only use it once. You multiply all the things they have in common (4) times all the things they don't have in common (3 and 7) SO... 4 * 3 * 7 = 84

OpenStudy (anonymous):

Mhmm........... Why would you multiply 3 and 7 even though they dont have them in common? I am so difficult :( I don't understand anything....

OpenStudy (anonymous):

You are ok... it is just the way you find the lcm. Do you have another one, I will help you.

OpenStudy (anonymous):

Yea 12 & 19 90*

OpenStudy (anonymous):

What is the 90* 12 = 3 * 4 19 = 1 * 19 so it is just 1 * 3 * 4 * 19 nothing in common which means it is actually 12*19

OpenStudy (anonymous):

I meant 90 not 19 lol

OpenStudy (anonymous):

I wasnt looking when i was typing

OpenStudy (anonymous):

OK 12 = 2 * 2 * 3 (4 times 3) 90 = 2 * 5 * 9 (2 times 45 and 45 is 5 times 9) Do you get this?

OpenStudy (anonymous):

Yea I understand a llittle...... Im struggling in math so I hope i'm not bothering.....

OpenStudy (anonymous):

OR 12 = 2 * 6 12 = 2 * 2 * 3 (I replaced the 6 with 2 * 3) 90 = 2 * 45 90 = 2 * 5 * 9 (I replaced the 45 with 5 * 9) 90 = 2 * 5 * 3 * 3 (I replaced the 9 with 3 * 3) This is how it should look.. I forgot to break down the 9 in my last post

OpenStudy (anonymous):

Oh now get it better!

OpenStudy (anonymous):

12 = 2 * 2 * 3 90 = 2 * 5 * 3 * 3 What do they both have....... answer (2 and a 3) So you multiply 2 * 3 (what they have in common) and then everything else 2 * 3 (common) * 2 * 5 * 3 (not common) 180

OpenStudy (anonymous):

Oh okay I get that

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

Anymore?

OpenStudy (anonymous):

Uhh 28, 42

OpenStudy (anonymous):

28x3=84 42x2=84. Your LCM is 84!